# Stabilizing Non-Abelian Topological Order against Heralded Noise via Local Lindbladian Dynamics

### v1

@@ V.3 Phase diagram of the $D_{4}$ model
We begin by discussing the correspondence between flag dynamics and the logical fidelity (41) shown in Fig. 11. When both $X$ and $Z$ flags are in the absorbing state, as noted above the densities of both plaquette and vertex defects closely track the respective flag densities and rapidly equilibrate to their steady-state values of $1/2$, indicating a fully decohered mixed-state. This proliferation of stabilizer defects leads to a rapid decay of the logical fidelities for both the $|\widetilde{0}\rangle_{L}$ and |+~$|\widetilde{+}\rangle_{L}$ states.
In the fully active phase, Fig. 11(c), where the density of both flags and stabilizer defects is small in the steady-state, we expect that the strings of errors that create stabilizer defects (which must lie along strings of flags) have a very high probability of being short. This is clearly borne out in our numerics, as seen in Fig. 11(c): the probability of decoding to the incorrect logical state is vanishingly small (i.e. the logical fidelity is essentially $1$ after decoding the steady-state, for all times shown). In other words, from the perspective of decodability, our fully active phase has non-Abelian $D_{4}$ topological order.
@@
which estimates the logical fidelity at time $t$ from the flag density. Here, $|n^{f}_{X,Z}=0\rangle$ is state where both $X$ and $Z$ flags on all sites are set to 00 and $\rho_{\text{flag}}$ is reduced density matrix of the flags. This upper bounds the full fidelity, since a state with no flags automatically has no defects or error strings, and thus will error-correct into the correct logical state. We verify in Fig. 13(b) that for $L\leq 30$, where we can perform the full stabilizer simulation, this is a good approximation for the full fidelity.
@@
As seen in Fig. 13 (c), the resulting fidelities are consistent with a decoding time that scales logarithmically with the system size, providing compelling evidence that our steady-state is in the same mixed-state phase as the pure $D_{4}$ topologically ordered state. Since correction can only remove flags from the boundaries of flag clusters, the timescale to recover the initial state should be determined by the size of the largest connected clusters of flags. In Appendix E, we show that this is indeed the case: we plot the distribution of flag-clusters as a function of their size, and find that the average size of the largest cluster increases logarithmically with increasing system size. We also expect that the partially active phase we have found here represents an intrinsically mixed-state TO (distinct from the imTO characterizing the active phases of the Toric code) since it encodes a non-trivial classical memory (see Refs. [40, 55] for a discussion of such intrinsically mixed TO phases).
@@
Figure 14: Determination of first order transition point based on the decay of flag densities towards the absorbing state in a finite-sized system: the time $\tau$ required for the erasure flag density to cross a finite constant value, chosen to be $n^{f}_{0}=0.65<1$, when initialized in a state with no flags, is plotted as a function of linear dimension $L$ for varying values of $X$ correction rates $\gamma_{x}$ close to the transition. The critical transition-rate is characterized by the value of $\gamma_{x}$ where the concavity of curves change from concave up (active phase for infinite system) to concave down (absorbing phase for infinite system). (a) The time $\tau$ for $n^{f}_{X}$ to reach $0.65$ as a function of system size for parameters near the transition from absorbing to partially X-active (where $Z$ flags are in their absorbing state) phase. Because the $X$ flags on a given honeycomb lattice evolve independently of the $Z$ flags (and the $X$ flags of the remaining two lattices), we only simulate $X$ flags of a single honeycomb lattice, allowing access to larger system sizes. The results are thus independent of $\gamma_{z}$. Each data point is obtained by averaging over $1000$ MC realizations. (b) $\tau$ associated with the $Z$ flag density is shown as a function of system size for $\gamma_{z}=500$ and varying values of $\gamma_{x}$ close to the transition from partially X-active to fully active phase. The data-points are obtained by averaging over 500 MC realizations.
Figure 15: Bistability at the first order transitions: the probability distribution $\rho$ of the density of number of flags in a Monte Carlo trajectory being $n_{f}$ is shown at a specific time $t_{*}$. The time $t_{*}$ is chosen such that the average density is fixed at arbitrary finite value ($n^{f}_{\text{avg}}=0.6$) for varying values of $\gamma_{x}$. The data is shown for $500-1000$ independent Monte Carlo realizations. (a-b) The transition from absorbing to partially active state is analyzed by monitoring the density of $X$ flags on one of the honeycomb lattices of the $D_{4}$ model. (a) Absorbing phase (shown for $\gamma_{x}=4$, independent of $\gamma_{z}$): the distribution has a single peak centered near the average value and it moves toward $n^{f}_{X}=1$. (b) Near the absorbing to partially active transition ($\gamma_{x}=6.45$, independent of $\gamma_{z}$): the distribution is bimodal where some of the Monte Carlo trajectories enter the absorbing state ($n^{f}_{X}=1$) but a finite number of trajectories remain active with a small density of $X$ flags, witnessed by a distinct secondary peak at $n^{f}_{X}<0.5$. In the inset, we demonstrate that the location of this secondary peak remains fixed for long times. Here the dashed lines show the time evolution of average $X$ flag density $\langle n^{f}_{X}\rangle_{\text{avg}}$. The solid dots show the average density of flags conditioned on Monte Carlo realizations that are still in the active state at that instant of time ($\approx$ area under the secondary peak). This shows that while the average density decays as a function of time, the conditioned density remains stable. The large scatter of conditioned-average for small $L$ at very late times results from most trajectories being in the absorbing state. (c-d) Similar results are shown for the transition from $X$-active to fully active at $\gamma_{z}=500$, controlled by the density of $Z$ flags. (c) Unimodal distribution of $Z$ flag density changes into (d) bimodal distribution at $\gamma_{x}=18.25$.
@@ V.4 First order transitions
Finally, we study the nature of the transitions separating these different steady-state phases, which we find to be first order. In the previous Section, we established that the active to absorbing state transition of the flags coincides precisely with the transition of the stabilizer system. As it is numerically less expensive to simulate the reduced dynamics of flags, here we exploit the aforementioned fact and present numerical results based on simulations of only the flag densities, allowing access to higher system-sizes.
@@
To study the nature of the transition, we begin by locating it precisely. Here, we will determine the location of the transition along the $\gamma_{z}=500$ cut in the phase diagram shown in Fig.1(b). For this, we exploit the fact that the time $\tau$ to reach a large flag density (in this case, $0.65$) grows exponentially in system size in the active phase, while in the absorbing phase it reaches an $L$-independent value for modest system sizes. We can thus use the change in the concavity of the $\tau$ vs $L$ curve on log-log scale to pinpoint the transition from the absorbing to the partially X-active phase. We note that the dynamics of $X$ flags on honeycomb lattice of a fixed color is independent of $X$ flags on the remaining lattices and all of the $Z$ flags. Hence, we locate this transition by only simulating the $X$ flags on a single honeycomb lattice which allows access to higher system sizes. In Fig. 14(a), we see that for $\eta=1,\gamma_{z}=500$, this happens at $\gamma_{x}=6.45\pm 0.03$ independent of the value of $\gamma_{z}$. Similarly, the transition from the partially active into the fully active state is observed at $\eta=1,\gamma_{z}=500,\gamma_{x}=18.25\pm 0.25$, as shown in Fig. 14(b).
@@
Having located the two transitions precisely, we now study the nature of the transition by monitoring the density of the $X$ and $Z$ flags as we cross each transition point. The probability distribution of the density of $X$ flags at the first transition point ($\gamma_{x}=6.45$) is shown in Fig. 15(b). The snapshot is taken at a time $t$ such that $\langle n^{f}_{X}(t)\rangle_{avg}=0.6$. We observe a bimodal structure: in a significant fraction of samples, the system has maintained a density of flags that is much smaller than $1$, leading to a peak at a low flag density $n^{f}_{X}\approx 0.2$; in the remaining samples, the system is already in the absorbing state, leading to a $\delta$-function peak at $n^{f}_{X}=1$. These two peaks are well separated in the space of flag densities, indicating that the probability of being in a state with an intermediate flag density is vanishingly small. Such bistable behavior is a hallmark of first-order transitions, where the order parameter changes discontinuously by shifting the relative weights of two distinct configurations [105, 106].
@@
The bistability is observed for a finite window of time, after which most of the trajectories are trapped in the absorbing state due to finite size effects. The time dependence of the average flag density shown in Fig. 14 suggests that this time-scale grows polynomially with $L$. To see that this is indeed the case, in the inset of Fig. 15(b) we show the density of $X$ flags computed by averaging only over those realizations that have not entered the absorbing state at a given time. This restricted density remains at a fixed value for a long time, indicating the robustness of the peak at low flag density $\langle n^{f}_{X}\rangle=n_{0}$. This demonstrates that (given the relatively narrow peak-widths) we can approximate the total density as a weighted average
$n^{f}_{X}(t)=P_{\text{absorb}}+n_{0}(1-P_{\text{absorb}})\,,$ (43)
@@
where $P_{\text{absorb}}$ is the fraction of trajectories that are in the fully occupied absorbing state. Thus the time-scale over which bistable behavior is observed (i.e. $P_{\text{absorb}}<1$) for a given system size is the same as $\tau$ estimated from the full densities in Fig. 14– that is, the time window grows as $poly(L)$ near the transition. In contrast, for $\gamma_{x}$ values deep in the active phase, the time to reach an absorbing configuration grows exponentially with system size.
@@
We note that this bimodality is sharply distinct from the time evolution seen for systems in the absorbing phase. In this case, at any given time the probability distribution is characterized by a single peak, which moves to the right (i.e. to higher $n^{f}_{X}$) with increasing time, until all the weight is concentrated on the absorbing state. Fig. 15(a) shows this for $\gamma_{x}=4$ (in the absorbing phase) evaluated at the time $t^{\prime}$ ($t^{\prime}\neq t$), for which the flag density $\langle n^{f}_{X}(t^{\prime})\rangle_{avg}=0.6$ is equal to that shown in Fig. 15(b).
@@
Figs. 15(c-d) shows similar data for the densities of $Z$ flags near the transition from the partially active phase (which is an absorbing state for the $Z$ flags) into the fully active phase. Again, bistable behavior is observed near the transition (panel (d)) while in the $Z$-absorbing phase, a single peak is observed about the mean flag density, which moves to the right as time increases until all the weight is in the $Z$ absorbing state $\langle n^{f}_{Z}\rangle=1$.
@@
Figure 16: $D_{4}$ model in the presence of unheralded noise: the logical error probability $1-F$ of system affected by a combination of heralded (rate $\phi_{e}$) and unheralded (rate $(1-\phi_{e})$) noise, along with the correction channel, is plotted as a function of time for varying values of $\phi_{e}$. Here, $F_{\psi_{0}}$ is the fidelity of the initial logical state $|\psi_{0}\rangle$ with the decoded state.. The noise is perfectly heralded when $\phi_{e}=1$ and the fraction of non-erasure errors increases with $\phi_{e}$. The data is shown for $L=18$ sized system, averaged over $250-500$ Monte Carlo realizations where all flags are initialized in the $n^{f}_{X_{j}}=n^{f}_{Z_{j}}=0$ state. The logical error probabilities (computed using the MWPM decoder) are shown for qubits initialized in (a) $|\widetilde{0}\rangle_{L}$ and (b) |+~$|\widetilde{+}\rangle_{L}$ logical states. (c) The time $\tau$ required for the fidelity to decrease below $1/2$ is plotted as a function of the unheralded noise fraction. The dashed lines are obtained by fitting the data points (except $\phi_{e}=0$) to a power-law $\tau=a(1-\phi_{e})^{b}$.
@@ V.5 Imperfect heralding
The specific nature of these unheralded noise processes depends on the details of the qubit platform. Here we phenomenologically model them as an uncorrelated depolarizing channel acting on qubits without raising any of the erasure flags. Let $\phi_{e}$ be the fraction of errors that are heralded by flags and converted into erasures (i.e., $\phi_{e}=1$ corresponds to the perfect heralding situation being considered so far). Then, the Lindblad superoperator for unheralded error is given by
@@
where $j$ runs over all qubits in the system. The heralded fraction of the noise is represented by the erasure noise model in Eq. (3), where we replace $\eta\rightarrow\eta\phi_{e}$. The time evolution of the state under the combined heralded and unheralded noise along with correction operations is generated by $\mathcal{L}_{\eta_{e}}+\mathcal{L}_{\eta_{p}}+\mathcal{L}_{C}$. The unheralded errors pair-create stabilizer defects that are not connected by strings of flags. Our correction protocol does not act on such defects if there are no flags in their vicinity. Hence, once such defects are created their evolution is effectively uncorrelated from that of their partner. In a long time limit, for any finite value of unheralded error rate, this results in a finite density of defects that are well-separated from their partners. We expect that upon pairing these defects using MWPM decoder, we will not get back to the initial state with certainty, meaning that our steady state has lost its topological properties. In Fig. 16(a)-(b), we demonstrate this by evaluating the fidelity of state obtained by decoding the time evolved state using MWPM decoder with respect to two different initial states in the $D_{4}$ ground-state space. The fidelity in the steady-state vanishes for all values of erasure fraction $\phi_{e}<1$.
@@
Figure 17: (a) The rate at which plaquette stabilizers $A_{p}$ are measured in the steady-state is shown for perfectly heralded dynamics in the $\gamma_{x}-\gamma_{z}$ parameter space of the $D_{4}$ model. We compute this rate in the Monte Carlo (MC) simulation of flags by counting the average number of $A_{p}$ measurements per plaquette proposed by the correction jump operators in a single MC sweep. The data is shown for an $L=48$ sized system at a time $t_{f}=20$ and averaged over $100$ MC realizations. The measurement rate goes to zero in the absorbing and partially active phase since all $Z$ flags are raised in this regime. The $Z$ correction protocol becomes inactive here. When the correction rates are sufficiently high, the system enters an active phase where both flag densities are low and hence the plaquette measurements are applied at a finite rate by the correction channel. (b) The measurement rate is shown for a fixed value of $Z$ correction rate $\gamma_{z}=500$ as a function of $X$ correction rate (horizontal cut along dashed line in (a)). The measurement rate jumps from zero to a finite value near the first order transition where the steady-state density of $Z$ flags changes discontinuously.

### v2

@@ V.3 Phase diagram of the $D_{4}$ model
We begin by discussing the correspondence between flag dynamics and the logical fidelity (42) shown in Fig. 11 (also see Fig. 1(b)). When both $X$ and $Z$ flags are in the absorbing state, as noted above, the densities of both plaquette and vertex defects closely track the respective flag densities and rapidly equilibrate to their steady-state values of $1/2$, indicating a fully decohered mixed-state. This proliferation of stabilizer defects leads to a rapid decay of the logical fidelities for both the $|\widetilde{0}\rangle_{L}$ and $|\widetilde{+}\rangle_{L}$ states.
In the fully active phase, Fig. 11(c), where the density of both flags and stabilizer defects is small in the steady-state, we expect that the strings of errors that create stabilizer defects (which must lie along strings of flags) have a very high probability of being short. This is clearly borne out in our numerics, as seen in Fig. 11(c): the probability of decoding to the incorrect logical state is vanishingly small (i.e., the logical fidelity is essentially $1$ after decoding the steady-state, for all times shown). In other words, from the perspective of decodability, our fully active phase has non-Abelian $D_{4}$ topological order.
@@
Figure 13: Two way connectivity in the $D_{4}$ active phase: We estimate the time required for noiseless dynamics $\eta=0$ to take the (quasi)-steady-state $\rho$ in the active phase ($\gamma_{x}=35,\gamma_{z}=500$ simulated upto $500$ MC sweeps) back to the pure $D_{4}$ ground state $|\psi_{0}\rangle$. To numerically access the larger system sizes, we estimate the fidelity of the time-evolved state based only on the flag densities using $F_{\text{flags}}$ (see Eq. (43)). (a) $1-F_{\text{flags}}(t)$ evaluated for various system sizes with linear dimension $L$. The data is obtained by averaging over $1000$ independent Monte Carlo realizations. (b) $1-F_{\text{flags}}(t)$ (blue) and $1-F_{\text{full}}(t)$ of the qubit system (orange) are plotted as a function of time for $L=30$. This demonstrates that $F_{\text{flags}}(t)<F_{\text{full}}(t)$ and hence recovery time-scales obtained using $F_{\text{flags}}$ upperbound the recovery time computed in terms of the full fidelity of the stabilizer system. (c) Time $\tau$ to achieve $1-F_{\text{flag}}(t)<\varepsilon$ is plotted as a function of system size $L$ for various cutoff values $\varepsilon$ on a linear-log scale plot. The dotted lines show best fits to the functional form $\alpha\log{(L/L_{0})}$. We note that the presence of even a single flag in a given trajectory leads to zero fidelity with $|n^{f}_{X,Z=0}\rangle$ state.
@@
which estimates the logical fidelity at time $t$ from the flag density. Here, $|n^{f}_{X,Z}=0\rangle$ is state where both $X$ and $Z$ flags on all sites are set to 0 and $\rho_{\text{flag}}$ is the reduced density matrix of the flags. This upper bounds the full fidelity, since a state with no flags automatically has no defects or error strings, and thus will error-correct into the correct logical state. We verify in Fig. 13(b) that for $L\leq 30$, where we can perform the full stabilizer simulation, this is a good approximation for the full fidelity.
@@
As seen in Fig. 13 (c), the resulting fidelities are consistent with a local correction time that scales logarithmically with the system size, providing compelling evidence that our steady-state is in the same mixed-state phase as the pure $D_{4}$ topologically ordered state. While the time-scale for the state to get close to the initial flag-free state in terms of fidelity increases logarithmically with increasing $L$, the local densities of flags approach zero in system-size independent time. Since correction can only remove flags from the boundaries of flag clusters, the timescale to recover the initial state should be determined by the size of the largest connected clusters of flags. In Appendix E, we show that this is indeed the case: we plot the distribution of flag-clusters as a function of their size, and find that the average size of the largest cluster increases logarithmically with increasing system size. We also expect that the partially active phase we have found here represents an intrinsically mixed-state TO (distinct from the imTO characterizing the active phases of the Toric code) since it encodes a non-trivial classical memory (see Refs. [41, 57] for a discussion of such intrinsically mixed TO phases).
@@
Figure 14: Determination of first order transition point based on the decay of flag densities towards the absorbing state in a finite-sized system: the time $\tau$ required for the erasure flag density to cross a finite constant value, chosen to be $n^{f}_{0}=0.65<1$, when initialized in a state with no flags, is plotted as a function of linear dimension $L$ for varying values of $X$ correction rates $\gamma_{x}$ close to the transition. The critical transition rate is characterized by the value of $\gamma_{x}$ where the curve changes from concave up (active phase for infinite system) to concave down (absorbing phase for infinite system). (a) The time $\tau$ for $n^{f}_{X}$ to reach $0.65$ as a function of system size for parameters near the transition from absorbing to partially X-active (where $Z$ flags are in their absorbing state) phase. Because the $X$ flags on a given honeycomb lattice evolve independently of the $Z$ flags (and the $X$ flags of the remaining two lattices), we only simulate $X$ flags of a single honeycomb lattice, allowing access to larger system sizes. Each data point is obtained by averaging over $1000$ MC realizations. (b) $\tau$ associated with the $Z$ flag density is shown as a function of system size for $\gamma_{z}=500$ and varying values of $\gamma_{x}$ close to the transition from partially X-active to fully active phase. The data-points are obtained by averaging over 500 MC realizations.
[…]
@@ V.4 First order transitions
Finally, we study the nature of the transitions separating these different steady-state phases, which we find to be first order. In the previous Section, we established that the active to absorbing state transition of the flags coincides precisely with the transition of the stabilizer system. As it is numerically less expensive to simulate the reduced dynamics of flags, here we exploit the aforementioned fact and present numerical results based on simulations of only the flag densities, allowing access to higher system sizes.
We emphasize that the flags are classical degrees of freedom and undergo a classical non-equilibrium absorbing state transition [81, 80, 108]. The density of active sites and the survival probability of the initial activity serve as order parameters for such transitions. The active phase corresponds to a non-vanishing probability of persistent activity out to times exponentially long in the system size, whereas the system becomes inactive after a finite time in the absorbing phase. The location of this transition can be determined from the time-scale after which activity disappears. In the present case, this can be determined from the time it takes for the density of active (i.e., unflagged) sites, given by $1-n^{f}_{X}$ or $1-n^{f}_{Z}$, to fall to small values – this changes from exponential in system-size in the active phase, to a constant value independent of system-size in the absorbing phase.
As in the equilibrium case, these transitions may exhibit a first-order or second-order character. However, unlike in the equilibrium case, there is no notion of free energy for such absorbing state transitions [108]. Nonetheless, we can still make an analogy to the equilibrium scenario by monitoring the behavior of order parameters. For instance, when the system undergoes a continuous transition (e.g., in the case of directed percolation), the spatial and temporal correlation lengths diverge at the critical point [81]. This results in the power-law decay of density of active sites at the critical point, which can be analyzed using standard scaling techniques analogous to equilibrium critical phenomena [80]. As we show in the following though, near the transition point of the models considered here, a significant fraction of Monte Carlo trajectories enter the absorbing state while the remaining trajectories maintain an active configuration. Such bistable behavior is instead more analogous to first-order equilibrium transitions [109]. Analysis of the discontinuous transitions to the absorbing state using the finite-size scaling methods generally involves modifications to the dynamics itself [110], which we do not pursue here. Instead, we observe that the system exhibits a bistability of both the absorbing and the active (with much lower flag density) configurations – also observed in various discontinuous absorbing phase transitions [111, 112, 80], and qualitatively resembling equilibrium first-order transitions.
@@
To study the nature of the transition, we begin by locating it precisely. Here, we will determine the location of the transition along the $\gamma_{z}=500$ cut in the phase diagram shown in Fig.1(b). For this, we exploit the fact that the time $\tau$ to reach a large flag density (in this case, $0.65$) grows exponentially in system size in the active phase, while in the absorbing phase it reaches an $L$-independent value for modest system sizes. We can thus use the change in the concavity of the $\tau$ vs $L$ curve on a log-log scale to pinpoint the transition from the absorbing to the partially X-active phase. We note that the dynamics of $X$ flags on the honeycomb lattice of a fixed color is independent of $X$ flags on the remaining lattices and all of the $Z$ flags. Hence, we locate this transition by only simulating the $X$ flags on a single honeycomb lattice, which allows access to higher system sizes. In Fig. 14(a), we see that for $\eta=1,\gamma_{z}=500$, this happens at $\gamma_{x}=6.45\pm 0.03$ independent of the value of $\gamma_{z}$. Similarly, the transition from the partially active into the fully active state is observed at $\eta=1,\gamma_{z}=500,\gamma_{x}=18.25\pm 0.25$, as shown in Fig. 14(b).
@@
Having located the two transitions precisely, we now study the nature of the transition by monitoring the density of the $X$ and $Z$ flags as we cross each transition point. The probability distribution of the density of $X$ flags at the first transition point ($\gamma_{x}=6.45$) is shown in Fig. 15(b). The snapshot is taken at a time $t$ such that $\langle n^{f}_{X}(t)\rangle_{avg}=0.6$. We observe a bimodal structure: in a significant fraction of samples, the system has maintained a density of flags that is much smaller than $1$, leading to a peak at a low flag density $n^{f}_{X}\approx 0.2$; in the remaining samples, the system is already in the absorbing state, leading to a $\delta$-function peak at $n^{f}_{X}=1$. These two peaks are well separated in the space of flag densities, indicating that the probability of being in a state with an intermediate flag density is vanishingly small. As discussed above, such bistable behavior is a hallmark of first-order transitions, where the order parameter changes discontinuously by shifting the relative weights of two distinct configurations [109, 110].
@@
The bistability is observed for a finite window of time, after which most of the trajectories are trapped in the absorbing state due to finite size effects. The time dependence of the average flag density shown in Fig. 14 suggests that this time-scale grows polynomially with $L$. To see that this is indeed the case, in the inset of Fig. 15(b) we show the density of $X$ flags computed by averaging only over those realizations that have not entered the absorbing state at a given time. This corresponds to the density of flags contained in the realizations near the secondary peak in Fig. 15(b). We observe that while the average density decays as a function of time, the conditioned density remains stable. The density among the surviving trajectories remains at a fixed value for a long time, indicating the robustness of the peak at low flag density $\langle n^{f}_{X}\rangle=n_{0}$. This demonstrates that (given the relatively narrow peak-widths) we can approximate the total density as a weighted average
$n^{f}_{X}(t)=P_{\text{absorb}}(t)+n_{0}(1-P_{\text{absorb}}(t))\,$, (44)
@@
where $P_{\text{absorb}}(t)$ is the fraction of trajectories that are in the fully occupied absorbing state. Thus, the time-scale over which bistable behavior is observed (i.e. $P_{\text{absorb}}<1$) for a given system size is the same as $\tau$ estimated from the full densities in Fig. 14– that is, the time window grows as $poly(L)$ near the transition. In contrast, for $\gamma_{x}$ values deep in the active phase, the time to reach an absorbing configuration grows exponentially with system size.
@@
We note that this bimodality is sharply distinct from the time evolution seen for systems in the absorbing phase. In this case, at any given time, the probability distribution is characterized by a single peak, which moves to the right (i.e., to higher $n^{f}_{X}$) with increasing time, until all the weight is concentrated on the absorbing state. Fig. 15(a) shows this for $\gamma_{x}=4$ (in the absorbing phase) evaluated at the time $t^{\prime}$ ($t^{\prime}\neq t$), for which the flag density $\langle n^{f}_{X}(t^{\prime})\rangle_{avg}=0.6$ is equal to that shown in Fig. 15(b).
@@
Figs. 15(c-d) show similar data for the densities of $Z$ flags near the transition from the partially active phase (which is an absorbing state for the $Z$ flags) into the fully active phase. Again, bistable behavior is observed near the transition (panel (d)) while in the $Z$-absorbing phase, a single peak is observed about the mean flag density, which moves to the right as time increases until all the weight is in the $Z$ absorbing state $\langle n^{f}_{Z}\rangle=1$.
@@
Figure 16: $D_{4}$ model in the presence of unheralded noise: the logical error probability 1F1-F of system affected by a combination of heralded (rate $\phi_{e}$) and unheralded (rate $(1-\phi_{e})$) noise, along with the correction channel, is plotted as a function of time for varying values of $\phi_{e}$. Here, $F_{\psi_{0}}$ is the fidelity of the initial logical state $|\psi_{0}\rangle$ with the decoded state. The data is shown for $L=18$ sized system, averaged over 250500250-500 Monte Carlo realizations where all flags are initialized in the $n^{f}_{X_{j}}=n^{f}_{Z_{j}}=0$ state. The logical error probabilities (computed using the MWPM decoder) are shown for qubits initialized in (a) $|\widetilde{0}\rangle_{L}$ and (b) $|\widetilde{+}\rangle_{L}$ logical states. (c) The time $\tau$ required for the fidelity to decrease below $1/2$ is plotted as a function of the unheralded noise fraction. The dashed lines are obtained by fitting the data points (except $\phi_{e}=0$) to a power-law $\tau=a(1-\phi_{e})^{b}$.
@@ V.5 Imperfect heralding
The specific nature of these unheralded noise processes depends on the details of the qubit platform. Here, we phenomenologically model them as an uncorrelated depolarizing channel acting on qubits without raising any of the erasure flags. Let $\phi_{e}$ be the fraction of errors that are heralded by flags and converted into erasures (i.e., $\phi_{e}=1$ corresponds to the perfect heralding situation being considered so far). Then, the Lindblad superoperator for unheralded error is given by
@@
where $j$ runs over all qubits in the system. The heralded fraction of the noise is represented by the erasure noise model in Eq. (6), where we replace $\eta\rightarrow\eta\phi_{e}$. The time evolution of the state under the combined heralded and unheralded noise, along with correction operations, is generated by $\mathcal{L}_{\eta_{e}}+\mathcal{L}_{\eta_{p}}+\mathcal{L}_{C}$. The unheralded errors pair-create stabilizer defects that are not connected by strings of flags. Our correction protocol does not act on such defects if there are no flags in their vicinity. Hence, once such defects are created their evolution is effectively uncorrelated from that of their partner. In a long time limit, for any finite value of unheralded error rate, this results in a finite density of defects that are well-separated from their partners. We expect that upon pairing these defects using MWPM decoder, we will not get back to the initial state with certainty, meaning that our steady state has lost its topological properties. In Fig. 16(a)-(b), we demonstrate this by evaluating the fidelity of the state obtained by decoding the time-evolved state using MWPM decoder with respect to two different initial states in the $D_{4}$ ground-state space. The fidelity in the steady-state vanishes for all values of erasure fraction $\phi_{e}<1$.
@@
Figure 17: (a) The rate at which plaquette stabilizers $A_{p}$ are measured in the steady-state is shown for perfectly heralded dynamics in the $\gamma_{x}-\gamma_{z}$ parameter space of the $D_{4}$ model. We compute this rate in the Monte Carlo (MC) simulation of flags by counting the average number of $A_{p}$ measurements per plaquette proposed by the correction jump operators in a single MC sweep. The data is shown for an $L=48$ sized system at a time $t_{f}=20$ and averaged over $100$ MC realizations. The measurement rate goes to zero in the absorbing and partially active phase (purple region) since all $Z$ flags are raised in this regime. In the active phase (yellow region), the plaquette measurements are applied at a finite rate by the correction channel. (b) The measurement rate is shown for a fixed value of $Z$ correction rate $\gamma_{z}=500$ as a function of $X$ correction rate (horizontal cut along dashed line in (a)). The measurement rate jumps from zero to a finite value near the transition where the steady-state density of $Z$ flags changes discontinuously.